# TIAM - A Metric for Evaluating Alignment in Text-to-Image Generation

### v1

@@ Abstract
The progress in the generation of synthetic images has made it crucial to assess their quality. While several metrics have been proposed to assess the rendering of images, it is crucial for Text-to-Image (T2I) models, which generate images based on a prompt, to consider additional aspects such as to which extent the generated image matches the important content of the prompt. Moreover, although the generated images usually result from a random starting point, the influence of this one is generally not considered. In this article, we propose a new metric based on prompt templates to study the alignment between the content specified in the prompt and the corresponding generated images. It allows us to better characterize the alignment in terms of the type of the specified objects, their number, and their color. We conducted a study on several recent T2I models about various aspects. An additional interesting result we obtained with our approach is that image quality can vary drastically depending on the latent noise used as a seed for the images. We also quantify the influence of the number of concepts in the prompt, their order as well as their (color) attributes. Finally, our method allows us to identify some latent seeds that produce better images than others, opening novel directions of research on this understudied topic.
@@ 1 Introduction
The ability to generate synthetic images with neural models made significant advancements from the advent of the first GANs [7, 20]. More recently diffusion-based models  [11, 24, 25, 17, 1] have further pushed the boundaries of image synthesis by progressively denoising an initial latent noise to generate high-quality images. In parallel to these progresses, the question of evaluating the quality of the generated synthetic images has always been a delicate issue and has become a research question in itself. To address this problem [28], several metrics were proposed [8, 26, 32] but they suffer from various limits [2, 4]. The most recent models are conditioned on textual image descriptions, allowing fine control of the output. It nevertheless added a challenge to evaluate their outputs, namely to estimate to which extent the synthetic image generated corresponds to the textual description it was conditioned on.
Firstly, although Text-to-Image (T2I) models demonstrate strong semantic and compositional capabilities, achieving a visually pleasing image that aligns with the desired condition often requires the generation of multiple images to obtain a suitable one. We propose that a reliable generative model should exhibit alignment with the condition specified in the prompt, irrespective of the initial latent noise. To address and study the variability of results, we introduce a novel metric to assess the success rate of generative models according to a prompt, Text-Image Alignment Metric [^1] Source code: https://github.com/grimalPaul/TIAM abbreviated as TIAM. The initial latent noise plays a crucial role in our metric, enabling us to investigate its impact. We have demonstrated that certain initial latent noise configurations outperform others. This finding suggests the possibility of selecting more effective initial latent noise to generate desired content.
Secondly, recent research efforts [5, 22, 25, 27] have shown that text-conditioned diffusion models suffer from three main issues relating to the text-image alignment, that is the alignment between the expected content expressed in the textual prompt and the content that is actually contained in the generated image: (i) catastrophic neglect where one or more elements of the prompt are not generated or sometimes mixed, (ii) attribute binding where attributes are bound to the wrong entities and (iii) attribute leaking where attributes specified in the prompt are correctly bound but some other elements in the scene are also wrongly bound with this attribute (Fig 1).
With TIAM, we propose to analyze the success rate of generative models under the scope of catastrophic and attribute-binding issues. For the latter, we propose a solid method to evaluate colors aligned with human perception.
@@
Figure 1: Respectively, pictures generated with the prompts “a photo of a lion and a bear”, “a photo of a blue cat and a yellow car”, and “a photo of a red bus driving down the street” generated with Stable diffusion v1.4. (a) The bear is missing, (b) the attributes are swapped (c) the wall becomes red because of the adjective that characterizes the bus.
Finally, the investigation of the influence of some words and attributes remains largely understudied. Tang et al. [27] provided some insights by examining the impact of words on generated outcomes. Using TIAM, we provide further insights into the relationship between textual conditioning and generative results. Relying on prompt template rather than natural prompts, TIAM allows quantifying the prompt-image alignment with regard to syntactic aspects, in particular the importance of the position of the main entities in the prompt.
In summary, our main contributions are the following: (i) We present a new metric based on prompt templates to quantify automatically the performances of T2I models in terms of prompt-image alignment (ii) We conduct an in-depth study of several diffusion models and quantify their behavior relating to catastrophic neglect and attribute binding issues (iii) We study the influence of the initial diffusion noise (seed) for all these models. The main insights resulting from this study are (a) the alignment performances of most T2I models drop significantly with the number of objects specified in the prompt (b) in practice there exists some seed that systematically leads to better results; TIAM allows us to identify them and they still provide better results with objects that are out of our study domain (c) most T2I models succeed in attributing color to one object but the performance drop even further with more objects.
@@ 2 Related Work
The evaluation of generated images often relies on metrics such as Inception Score (IS) [26] and Fréchet Inception Distance (FID) [8], which are commonly used to assess their quality and fidelity. However, these metrics do not capture the alignment between the provided condition(s) and the generated images. To address this issue, several methods have been proposed to measure whether the content of the generated images reflects the given conditions. One approach involves utilizing classification models [23] and object detection models (Semantic Object Accuracy, abbreviated as SOA) [10] to determine whether the generated images contain the specified objects or adhere to certain criteria. SOA uses actual prompts from COCO while our approach relies on prompt templates that allows a finer analysis of the influence of each element in the prompt. Another method specific to text-image alignment involves employing visual score similarity measures. It leverages contrastive models such as CLIP, which compute a similarity score between representations of images and text. Although CLIP has a great average semantic representation, it has poor compositional understanding [31], limiting a fine evaluation of text-image alignment. Our metric addresses this issue by leveraging a high-performance detector and controlling the requested attributes in an explainable manner and aligning with human perception.
Some benchmarks such as DrawBench [25] propose a limited set of text prompts covering 11 skills to evaluate models. To gain a more comprehensive understanding of the performance of T2I models, we approach the issue by focusing on catastrophic neglect and attribute binding. To overcome results that are very sensitive to asked objects in prompt and evaluate the model’s performance accurately, we propose a methodology that tests all combinations of objects and attributes and generates multiple images. By testing all the combinations in the prompts, we can assess the model’s success rate in capturing and representing the desired attributes. This approach offers a comprehensive evaluation, free from the limitations of specific labels, resulting in a more precise and less biased assessment.
@@ 3 Method
We propose a new method to measure the success rate of generative models covering catastrophic neglect and/or attribute binding. First, we generate multiple prompts given a set of word labels and possibly attributes. We then generate multiple images for each prompt and detect if the expected asked elements are detected on the image. Finally, we compute the final score.
@@ 3.1 TIAM Text-Image Alignment Metric
Our approach is based on templates that are used to generate sets of prompts. We adopt a formalism inspired by the disentangled representation theory [9, 29]. The prompts contain $N$ objects, each of which can be qualified by an attribute. Hence the object at position $i$ is a token belonging to the set $\mathcal{O}_{i}$ and qualified by an attribute in the set $\mathcal{A}_{i}$. For example, let us consider the template “a photo of det($o_{1}$, $a_{1}$) $a_{1}$ $o_{1}$ and det($o_{2}$, $a_{2}$) $a_{2}$ $o_{2}$” where $o_{i}\in\mathcal{O}_{i}$, $a_{i}\in\mathcal{A}_{i}$, and det($o_{i}$, $a_{i}$) is a determinant that depends on the object or attribute if present. If $\mathcal{O}_{1}=\mathcal{O}_{2}=${‘car’,‘bike’,‘truck’} and $\mathcal{A}_{1}=\mathcal{A}_{2}=${‘blue’,‘green’,‘red’}, then it can produce prompts such as “a photo of a blue car and a red truck” or “a photo of a green bike and a blue car”.
@@
where the prompt is instanced from the template $t$ and its “latent concepts” $z\in\mathcal{Z}$, $y(z)$ is the labels that relate to the expected content of the synthetic image generated by the model $G()$ (conditioned by the prompt) from a seed (using to generate a latent noise) $\chi$. $f$ is a scoring function that compares the ground truth $y(z)$ to the output of a model that detects objects or produces segmentation maps from the synthetic image, the resulting score being in $\{0,1\}$ depending on whether the content matches the ground truth or not (see subsection 3.3).
With the definition of Equation 1, a template can generate $\sum_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$ different prompts, where |.||.|| . | is the cardinal of the set. However, we restrict in practice the inference of the template such that an object or an attribute can not appear twice in the prompt. In a simple case where the attributes and the objects set are the same at each position, the number of unique prompts is given by:
Proposition 1: if $|\mathcal{O}|\geq N$, $|\mathcal{A}|\geq N$ and $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$ and $\forall(i,j)\in[\![1,N]\!]^{2}$ s.t $i<j$, we force $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$, thus the number of unique prompts generated in the context of Equation 1 is $\frac{|\mathcal{O}|!|\mathcal{A}|!}{\left(|\mathcal{O}|-N\right)!\left(|\mathcal{A}|-N\right)!}$
The proof is given in the supplementary material A. We also derive the more general case with possibly different attributes and objects at each position. The global evaluation process is summarized in Fig 2.
Figure 2: Overview of the evaluation pipeline. (1) Generate a dataset of prompts (2) Generate $n\geq 16$ images per prompt (3) Detect if the requested labels are present in the image (4) Compute TIAM. In this example, we do not define attributes.
@@ 3.2 Attribute definition
We study the color attributes in this paper, but this metric could be applied under the scope of another type of attribute (e.g., size or texture). Selecting colors that are aligned with human perception is not trivial because of the infinity of possibilities. We base our choice on the work of Berlin and Kay [3]. They define eleven universal basic colors $\mathcal{C}$: white, black, red, green, blue, brown, purple, pink, orange, yellow, and grey. They asked individuals to select an array of 329 colors (provided by the Munsell Colors Company), the chips that correspond to each basic color, and to select the most typical examples. We use the results of the most typical examples of American English and convert the colors from the Munsell System to the CIELAB space. This one offers the advantage of a color distribution more aligned with human perception and a suitable space to compare color differences.
We remark that brown and orange were respectively too close to black and white and we remove them (more information provided in supplementary material C. We set
@@
However, we use the following set as reference colors (to compare with $\mathcal{A}$): white, black, red, green, blue, purple, pink, yellow.
To define if $\mathcal{G()}$ correctly assigns $a_{i}$ and $o_{i}$, we use a segmentation model that delimits $o_{i}$ if present on the image. We compare all the pixels inside the segmentation maps with our reference colors in the CIELAB space. If we detect at least a proportion of 40% of $a_{i}$, we consider a successful binding.
@@ 3.3 Main implementation details
The template $t$ always includes the start mention “a photo of”. Indeed, in our case, the detector is trained on real images and we suppose that it would therefore have more difficulty managing e.g. sketches or 3D renders. The general form of the templates is that given as an example above, by varying the syntactic context. For the detection and segmentation tasks we use YOLOv8 [13], a state-of-the-art model pre-trained on the 80 COCO labels [15]. We fixed the object confidence threshold to 0.25 but conducted a study asserting that a higher threshold (0.4 to 0.8) does not change the relative order performance of the models. The results can be found in the supplementary material F, where we provide a detailed description of the labels used in each experiment and their combination within the generated prompts.
The scoring function $f$ mainly focuses on true positives, considering the alignment as successful if we find at least one of each named object in the prompt. If the objects are characterized by an attribute, each object must be present with its correct attribute at least once in the image to be marked as a success. The criterion is therefore strict on the alignment aspect, requiring the presence of all the named objects and attributes. This approach differs from CLIP scores, which provide a general trend without reflecting a comprehensive explanation of the results. It nevertheless does not penalize the presence of other objects (false positive) since it is not prohibitive in a generative context. $f$ prevent also element mixing. Indeed, if the segmentation masks of two different objects overlap with an IoU value of 0.95 or higher, we remove both objects before calculating the score with the remaining detected objects.
@@ 4 Study
We conduct an in-depth study to characterize and quantify the limits of several diffusion models, which differ by the text-conditioning method, the architecture, or the inference process. We consider Stable Diffusion v1-4 (SD 1.4) [^2] https://huggingface.co/CompVis/stable-diffusion-v1-4 and Stable Diffusion v2 (SD 2) [^3] https://huggingface.co/stabilityai/stable-diffusion-2-base based on the LDM model [24] that uses a fixed pretrained text encoder CLIP [19], which is based on an auto-regressive architecture. Both versions are trained with the same variational autoencoder (VAE), but two different U-nets (the part that learns to denoise image). The CLIP models that guide the U-nets also differ between SD 1.4 and SD 2. In addition, we evaluate the two SD models with Attend-and-Excite (A&E) [5], an optimized inference process using the cross-attention maps to attend the subject tokens in the prompt. We follow the author’s implementation and only attend to the token of object $o_{i}$, even if an attribute characterizes the object. We also consider an unCLIP model [22, 14] conditioned with a CLIP image prior and a CLIP text embedding. Finally, we study DeepFloyd IF (IF) [^4] https://github.com/deep-floyd/IF a cascade diffusion inspired by [25] conditioned using a T5 XXL [21] text encoder. Other implementation details are reported in the supplementary material G.

### v2

@@ Abstract
The progress in the generation of synthetic images has made it crucial to assess their quality. While several metrics have been proposed to assess the rendering of images, it is crucial for Text-to-Image (T2I) models, which generate images based on a prompt, to consider additional aspects such as to which extent the generated image matches the important content of the prompt. Moreover, although the generated images usually result from a random starting point, the influence of this one is generally not considered. In this article, we propose a new metric based on prompt templates to study the alignment between the content specified in the prompt and the corresponding generated images. It allows us to better characterize the alignment in terms of the type of the specified objects, their number, and their color. We conducted a study on several recent T2I models about various aspects. An additional interesting result we obtained with our approach is that image quality can vary drastically depending on the noise used as a seed for the images. We also quantify the influence of the number of concepts in the prompt, their order as well as their (color) attributes. Finally, our method allows us to identify some seeds that produce better images than others, opening novel directions of research on this understudied topic.
@@ 1 Introduction
The ability to generate synthetic images with neural models made significant advancements from the advent of the first GANs [15, 36]. More recently diffusion-based models  [20, 40, 41, 33, 1] have further pushed the boundaries of image synthesis by progressively denoising an initial noise to generate high-quality images. In parallel to these advances, the question of evaluating the quality of these synthetic images has always been a delicate issue and has become a research question in itself. To address this problem [45], several metrics were proposed [17, 42, 50] but they suffer from various limits [2, 5]. The most recent models are conditioned on textual image descriptions, allowing fine control of the output. It nevertheless adds a challenge to evaluate their outputs, namely to estimate to which extent the synthetic image generated corresponds to the textual description it was conditioned on.
Although Text-to-Image (T2I) models demonstrate strong semantic and compositional capabilities, achieving a visually pleasing image that aligns with the desired condition often requires the generation of multiple images to obtain a suitable one. A reliable generative model should exhibit alignment with the condition specified in the prompt, irrespective of the starting noise. To address and study the variability of results, we introduce a novel metric to assess the success rate of generative models according to a prompt, Text-Image Alignment Metric [^1] Source code: https://github.com/grimalPaul/TIAM (TIAM). The initial noise plays a crucial role in our metric, enabling us to investigate its impact. We show that certain initial noise configurations outperform others, suggesting the possibility of selecting them to get better synthetic images.
Recent research efforts [6, 38, 41, 44, 11] have shown that text-conditioned diffusion models suffer from three main issues related to the alignment between the expected content expressed in the textual prompt and the one actually generated in the image: (i) catastrophic neglect, where one or more elements described in the prompt are not generated or sometimes mixed, (ii) attribute binding, where attributes (e.g. color) are bound to the wrong entities, and (iii) attribute leaking, where attributes specified in the prompt are correctly bound but some other elements in the scene are also wrongly bound with this attribute (Fig. 1).
With TIAM, we propose to analyze the success rate of generative models under the scope of catastrophic and attribute-binding issues. For the latter, we propose a solid method to evaluate color alignment with human perception.
@@
Figure 1: Images generated with the prompts “a photo of a lion and a bear”, “a photo of a blue cat and a yellow car”, and “a photo of a red bus driving down the street” generated with Stable diffusion v1.4. (a) The bear is missing, (b) the attributes are swapped, (c) the bus color (red) leaks on the wall
To date, the investigation of the influence of some words and attributes remains largely understudied. Tang et al. [44] provide some insights by examining the impact of words on generated outcomes. Using TIAM, we provide further insights into the relationship between textual conditioning and generative results. Relying on prompt templates rather than natural prompts, TIAM allows quantifying the prompt-image alignment w.r.t syntactic aspects, in particular the importance of the position of the main entities in the prompt.
In summary, our main contributions are: (i) a new metric based on prompt templates to quantify automatically the performance of T2I models in terms of prompt-image alignment; (ii) an in-depth study of several diffusion models and quantification of their behavior relating to catastrophic neglect and attribute binding issues; (iii) a study on the influence of the initial diffusion noise (seed) for all these models. The main insights resulting from this study are (a) the alignment performance of most T2I models drops significantly with the number of objects specified in the prompt; (b) in practice there exists some seeds that systematically lead to better results; TIAM allows us to identify them and they still provide better results with objects that are out of our study domain; (c) most T2I models succeed in attributing color to one object but the performance drop with more objects.
@@ 2 Related Work
The evaluation of generated images often relies on metrics such as Inception Score (IS) [42] and Fréchet Inception Distance (FID) [17], which are commonly used to assess their quality and fidelity. However, these metrics do not capture the alignment between the provided condition(s) and the generated images. To address this issue, several methods have been proposed to measure whether the content of the generated images reflects the given conditions. One approach involves utilizing classification models [39] and object detection models (Semantic Object Accuracy, abbreviated as SOA) [19] to determine whether the generated images contain the specified objects or adhere to certain criteria. SOA uses actual prompts from COCO while our approach relies on prompt templates that allow a finer analysis of the influence of each element in the prompt. Another method specific to text-image alignment involves employing visual score similarity measures. It leverages contrastive models such as CLIP, which compute a similarity score between representations of images and text. Although CLIP has a great average semantic representation, it has poor compositional understanding [49], limiting a fine evaluation of text-image alignment. Our metric addresses this issue by leveraging a high-performance detector and by controlling the requested attributes in an explainable manner and aligning with human perception.
Recent efforts have introduced innovative methodologies to assess skills and measure biases of T2I models. Notably, Drawbench [41] presents a limited set of text prompts covering 11 skills to evaluate the models. Similarly, DALL-EVAL [7] proposes to evaluate three visual reasoning skills: the capacity to generate one object (object recognition), the capacity to generate the exact quantity number of the asked object (counting), and the capacity to place two objects (spatial reasoning). Additionally, they probe gender and skin bias in the model representation. Another contribution by Zhang et al. [51] delves into the gender depiction disparities enabling the study of potential stereotypes. Our methods focus on catastrophic neglect and attribute binding. We can see the catastrophic neglect approach as object recognition but by testing one or more different objects in the conditioning prompt. Moreover, to overcome results that are very sensitive to asked objects in prompt and evaluate the model’s performance accurately systematically, we explore all various object and attribute combinations and we generate multiple images per prompt to enhance result robustness. Recognizing the impact of initial noises, we advocate for multiple seed testing. This approach offers a comprehensive evaluation, free from the limitations of specific labels, resulting in a more precise and less biased assessment. A closely related work [14] conducted concurrently with ours shares some conclusions, though it does not delve into the aspect of the starting noises.
@@ 3 Method
We propose a new method to measure the success rate of generative models w.r.t catastrophic neglect and/or attribute binding. First, we generate multiple prompts given a set of word labels and possible attributes. We then generate multiple images for each prompt and detect if the expected elements are present on the image, leading to the final score.
@@ 3.1 TIAM Text-Image Alignment Metric
Our approach is based on templates that are used to generate sets of prompts. We adopt a formalism inspired by the disentangled representation theory [18, 46]. The prompts contain $N$ objects, each of which can be qualified by an attribute. Hence the object at position $i$ in a prompt is a token belonging to the set $\mathcal{O}_{i}$ and qualified by an attribute in the set $\mathcal{A}_{i}$. For example, let us consider the template “a photo of det($o_{1}$, $a_{1}$) $a_{1}$ $o_{1}$ and det($o_{2}$, $a_{2}$) $a_{2}$ $o_{2}$” where $o_{i}\in\mathcal{O}_{i}$, $a_{i}\in\mathcal{A}_{i}$, and det($o_{i}$, $a_{i}$) is a determinant that depends on the object or attribute if present. If $\mathcal{O}_{1}=\mathcal{O}_{2}=${‘car’,‘bike’,‘truck’} and $\mathcal{A}_{1}=\mathcal{A}_{2}=${‘blue’,‘green’,‘red’}, then it can produce prompts such as “a photo of a blue car and a red truck” or “a photo of a green bike and a blue car”.
@@
where the prompt is instanced from the template $t$ and its “latent concepts” $z\in\mathcal{Z}$, $y(z)$ is the labels that relate to the expected content of the synthetic image generated by the model $G()$ (conditioned by the prompt) from a seed (using to generate an initial noise) $\chi$. $f$ is a scoring function that compares the ground truth $y(z)$ to the output of a model that detects objects or produces segmentation maps from the synthetic image, the resulting score being in $\{0,1\}$, depending on whether the content matches the ground truth or not (see Section 3.3).
With the definition of Equation 1, a template can generate $\sum_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$ different prompts, where |.||.|| . | is the cardinal of the set. However, we restrict in practice the inference of the template such that an object or an attribute can not appear twice in the prompt. In a simple case where the attribute and object sets are the same at each position, the number of unique prompts is given by:
Proposition 1:
if $|\mathcal{O}|\geq N$, $|\mathcal{A}|\geq N$, and $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$, and $\forall(i,j)\in[\![1,N]\!]^{2}$ s.t $i<j$, we force $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$, thus the number of unique prompts generated in the context of Equation 1 is $\frac{|\mathcal{O}|!|\mathcal{A}|!}{\left(|\mathcal{O}|-N\right)!\left(|\mathcal{A}|-N\right)!}$
The proof is given in the Supp. Mat. A (supplementary material). We also derive a more general case with possibly different attributes and objects at each position. The global evaluation process is summarized in Fig. 2.
Figure 2: Overview of the evaluation pipeline. (1) Generate a dataset of prompts. (2) Generate $n\geq 16$ images per prompt. (3) Detect if the requested labels are present in the images. (4) Compute TIAM. In this example, we do not define attributes.
@@ 3.2 Attribute definition
To define if $G()$ correctly assigns $a_{i}$ and $o_{i}$, we use a segmentation model that delimits $o_{i}$ if present on the image. We compare all the pixels inside the segmentation maps with our reference colors in the CIELAB space. If we detect at least a proportion of 40% of $a_{i}$, we consider a successful binding.
@@ 3.3 Main implementation details
The template $t$ always includes the start mention “a photo of”. Indeed, in our case, the detector is trained on real images and we suppose that it would therefore have more difficulty managing e.g. sketches or 3D renders. The general form of the templates is that given as an example above, by varying the syntactic context.
For the detection and segmentation tasks, we use YOLOv8 [23], a state-of-the-art model pre-trained on the 80 COCO labels [31]. We set the confidence threshold to 0.25 but conducted a study asserting that a higher threshold (0.4 to 0.8) does not change the relative order performance of the models. The results can be found in the Supp. Mat. F, with a detailed description of the labels used in each experiment and their combination within the generated prompts.
The scoring function $f$ mainly focuses on true positives, considering the alignment as successful if we find at least one of each named object in the prompt. If the objects are characterized by an attribute, each object must be present with its correct attribute at least once in the image to be marked as a success. The criterion is therefore strict on the alignment aspect, requiring the presence of all the named objects and attributes. This approach differs from CLIP scores, which provide a general trend without reflecting a comprehensible explanation of the results. It nevertheless does not penalize the presence of other objects (false positive) since it is not prohibitive in a generative context. Indeed, if the segmentation masks of two different objects overlap with an IoU value of 0.95 or higher, we remove both objects before calculating the score with the remaining detected objects.
@@ 4 Study
We conduct an in-depth study to characterize and quantify the limits of several diffusion models, which differ by the text-conditioning method, the architecture, or the inference process. We consider Stable Diffusion v1-4 (SD 1.4) and Stable Diffusion v2 (SD 2) based on the LDM model [40] that uses a fixed pre-trained text encoder CLIP [35], which is based on an auto-regressive architecture. Both versions are trained with the same variational autoencoder (VAE), but two different U-nets (the part that learns to denoise image). The CLIP models that guide the U-nets also differ between SD 1.4 and SD 2. In addition, we evaluate the two SD models with Attend-and-Excite (A&E) [6], an optimized inference process that uses the cross-attention maps to attend the subject tokens in the prompt. We follow the authors’ implementation and only attend to the token of object $o_{i}$, even if an attribute characterizes the object. We also consider an unCLIP model [38, 27] conditioned with a CLIP image prior and a CLIP text embedding. Finally, we study DeepFloyd IF (IF), a cascade diffusion inspired by [41], conditioned using a T5 XXL [37] text encoder. Other details are reported in the Supp. Mat. G.